Raising rabbits. Before Roland sold two female rabbits, half of his rabbits were female. After the sale, only one-third of his rabbits were female. If x represents his original number of rabbits, then
1/2x-2=1/3(x-2)

Answer by ptaylor(2198)

x=original number of rabbits
Then x/2=original number of female rabbits
and x/2=original number of male rabbits

After he sells two female rabbits, he has 1/3(x-2) female rabbits left
So, the way I see it is as follows:
Number of female rabbits after the sale {(1/3)(x-2)} plus the number of male rabbits (x/2) equals the original number of rabbits less 2 (the females that he sold):
1/3(x-2)+(x/2)=x-2 simplify
(1/3)x-(2/3)+(x/2)=x-2 multiply each term by 6
2x-4+3x=6x-12 simplify
5x-4=6x-12 subtract 6x from and add 4 to both sides
5x-6x-4+4=6x-6x-12+4
-x=-8 multiply each term by -1
x=8 original number of rabbits
CK
Original number of females=x/2=8/2=4
Give away 2, then Roland has 2 females left and 4 males left for a total of 6 rabbits and 1/3 of them, or 2, are females.

Hope this helps---ptaylor